Mike is 2 years younger than three times Jill’s age. The sum of their ages is 38. Which equation can be used to solve for Mike and Jill’s age if x represents how old Jill is?

Found 2 solutions by mananth, rothauserc:
Answer by mananth(16946)

Mike is 2 years younger than three times Jill’s age. The sum of their ages is 38. Which equation can be used to solve for Mike and Jill’s age if x represents how old Jill is?
jill is x years
Mike is 3x-2 years
sum of their ages is 38
x+ 3x-2 =38

Answer by rothauserc(4718)

let x be Jill's age and m be Mike's age
:
1) m = 3x -2
:
2) m + x = 38
:
using equation 1, substitute for m in equation 2
:
3x -2 +x = 38
: